### v2

# Transport alpha divergences

Abstract.
We derive a class of divergences measuring the difference between probability density functions on a one-dimensional sample space. This divergence is a one-parameter variation of the Itakura–Saito divergence between quantile density functions. We prove that the proposed divergence is one-parameter variation of transport Kullback-Leibler divergence and Hessian distance of negative Boltzmann entropy with respect to Wasserstein-$2$ metric. From Taylor expansions, we also formulate the $3$-symmetric tensor in Wasserstein space, which is given by an iterative Gamma three operators. The alpha-geodesic on Wasserstein space is also derived. From these properties, we name the proposed information measures transport alpha divergences. We provide several examples of transport alpha divergences for generative models in machine learning applications.
Key words and phrases:
Transport alpha divergence; Quantile density function; Transport Hessian metric; Transport 3–symmetric tensor; Gamma 33 calculus.
1. Introduction
Information measures of differences between probability distributions play essential roles in statistics, information theory, signal processing, and estimation [1, 7, 10, 11]. It is a generalization of Kullback-Leibler (KL) divergence, with dualities and variational properties. One typical example is the Alpha divergence, which has vast applications in machine learning inference problems and Bayesian sampling problems.
Information geometry (IG) studies the geometric, duality, and invariance properties of divergences in probability density space. Examples include Kullback-Leibler (KL) divergence, alpha-divergences, and their generalizations [1, 7, 10, 31]. In this study, the derivatives of negative Boltzmann–Shannon entropy in $L^{2}$ space are with fundamental roles. Its first-order derivative is the likelihood function, its second derivative forms the Fisher-Rao metric, while its third derivative introduces the Amari–Chenstov tensor. The combination of these derivatives characterizes the KL divergence and its one-parameter variation, namely alpha divergences, where alpha is a scalar. From these characterizations, IG studies and constructs finite-dimensional probability models, with desirable approximation and convexity properties in inference problems.
Recently, optimal transport studies the other distance function, namely Wasserstein distances, in probability density space [30]. The Wasserstein distance introduces the dualities based on pushforward mapping functions, which nowadays have vast applications in estimation and AI sampling problems, such as generative adversarial networks [4]. In particular, there also exists a Riemannian metric for Wasserstein-$2$ distance in probability density space [14, 28, 30]. Under this Wasserstein-$2$ metric, the derivatives of Boltzman-Shannon entropy are of importance in simulating physics equations [23, 13] and Ricci curvature lower bound in a sample space [8, 22, 25]. The study of first- and second-order derivatives in Wasserstein-$2$ space has also been used in statistics and optimization in machine learning algorithms [15]. A natural question arises. What is the alpha divergence in terms of pushforward mapping functions?
This paper answers this question by applying information geometry methods to optimal transport geometry. For simplicity of presentation, we focus on the result in one-dimensional sample space. We show that the transport alpha divergence is a one-parameter family, which interpolates the transport KL divergence function [17] and transport Hessian distances [18]. We derive the third order derivative, a.k.a. 33–symmetric tensor, of negative Boltzmann–Shannon entropy in Wasserstein-$2$ space. Several properties of transport alpha divergences are presented, including duality relation, Taylor expansions, generalized Bregman divergences, and generalized Pythagorean theorem in Wasserstein-$2$ spaces.
We briefly present the main result. Given a one-dimensional domain $\Omega$ and two strictly positive probability density functions $p$, $q$, we propose the alpha divergence in Wasserstein-$2$ space as
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\int_{0}^{1}\Big{(}\big{(}\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}\big{)}^{\alpha}-{\alpha}\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}-1\Big{)}dx,&\qquad\textrm{if $\alpha\neq 0$;}\\ &\frac{1}{2}\int_{0}^{1}|\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}|^{2}q(x)dx,&\qquad\textrm{if $\alpha=0$.}\end{aligned}\right$. (1)
where $Q_{p}$, $Q_{q}$ are quantile functions of $p$, $q$, respectively, and $Q^{\prime}_{p}$, $Q^{\prime}_{q}$ are derivatives of quantile functions, namely quantile density functions. We note that the quantile function is the inverse function of cumulative distribution function. We remark that compared with alpha divergences in $L^{2}$ space, the transport alpha divergence reformulates density functions to quantile density functions.
In literature, several joint studies exist among information geometry, optimal transport, and alpha divergences [5, 21, 24, 27]. For example, [24, 27] studies the optimal transport over a Bregman divergence ground cost. [21] studies the matrix decomposition viewpoint for different information metrics on Gaussian families. Recently, [5] also studies Wasserstein-2 metric with general Riemannian ground metric, and then study canniocal divergences from Wasserstein-2 distances. Compared to the above studies, we focus on the Hessian metric of negative Boltzman-Shannon entropies in Wasserstein-$2$ space. In this paper, we apply Hessian structure [1, 26] to construct divergence functionals in terms of Jacobi operators of mapping functions; see related works in [15, 16, 17, 18].
This paper is organized as follows. In section 2, we briefly review the definition of classical alpha divergence in positive octant and its relation with information geometry methods. In section 4, we first construct transport alpha divergence in one-dimensional sample space. Its Taylor expansions show both the Hessian metric and the 3-symmetric tensor in Wasserstein-$2$ space. Properties of transport alpha divergences, including generalized Bregman divergences and Pythagorean theorem in Wasserstein-$2$ spaces, are discussed. Several analytical formulas in generative models are provided in section 5.
2. Divergence functions and Information geometry methods
In this section, we briefly review alpha divergence functions in positive octant. We also recall information geometry methods for studying these divergences functionals [1, 2, 7].
From now on, we mainly follow derivations in [7, section 2.5.2]. Denote a dd–dimensional positive octant by $(0,+\infty)^{d}$. For any vectors $m=(m_{i})_{i=1}^{d}$, $n=(n_{i})_{i=1}^{d}\in(0,+\infty)^{d}$, the alpha divergence is defined by
$\mathrm{D}_{\alpha}[m\|n]=\sum_{i=1}^{d}f_{\alpha}(\frac{m_{i}}{n_{i}})n_{i},$
where $f_{\alpha}\colon\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is a function parameterized by a scalar $\alpha$, such that
$f_{\alpha}(z)=\left\{\begin{aligned} \frac{4}{1-\alpha^{2}}\Big{(}\frac{1-\alpha}{2}+\frac{1+\alpha}{2}z-z^{\frac{1+\alpha}{2}}\Big{)},&\quad\alpha\neq\pm 1;\\ z\log z-(z-1),&\quad\alpha=1;\\ -\log z+(z-1),&\quad\alpha=-1.\end{aligned}\right$.
Here $\log$ is the natural logarithm function. The alpha divergence is a distance-like function, namely divergence or contrast function that satisfies the following properties.
$\mathrm{D}_{\alpha}[m\|n]\geq 0;\quad\mathrm{D}_{\alpha}[m\|n]=0,\quad\textrm{iff}\quad m=n.$
We note that, in general when $\alpha\neq 0$, alpha divergence is not a distance function, since $\mathrm{D}_{\alpha}(m|n)\neq\mathrm{D}_{\alpha}(n|m)$. The following dual relation holds
$\mathrm{D}_{\alpha}[m\|n]=\mathrm{D}_{-\alpha}[n\|m].$
There are three important examples of alpha divergences, widely used in statistical inference applications.
  • (i)
    $\alpha=0$: Squared Hellinger distance (up to a scaling factor)
    $\mathrm{D}_{0}[m\|n]=2\sum_{i=1}^{d}(\sqrt{m_{i}}-\sqrt{n_{i}})^{2}.$
  • (ii)
    $\alpha=1$: Kullback-Leibler (KL) divergence
    $\mathrm{D}_{1}[m\|n]=\sum_{i=1}^{d}m_{i}\log\frac{m_{i}}{n_{i}}-(m_{i}-n_{i}).$
  • (iii)
    $\alpha=3$: Chi-squared divergence
    $\mathrm{D}_{3}[m\|n]=\frac{1}{2}\sum_{i=1}^{d}\frac{(m_{i}-n_{i})^{2}}{n_{i}}.$
The alpha divergence has several important properties from Hessian structures of an entropy function, especially Taylor expansions and alpha geodesics. Denote a finite dimensional Boltzman-Shannon entropy function by $H(m)=-\sum_{i=1}^{n}m_{i}\log m_{i}$. Denote the Hessian matrix of negative $H$, also named Fisher matrix, by
$g_{ij}(m)=-\frac{\partial^{2}}{\partial m_{i}\partial m_{j}}H(m)=\frac{1}{m_{i}}\delta_{ij},\quad\textrm{for $i,j\in\{1,\cdots,d\}$;}$
and denote the third derivative of $H$ by a 3–symmetric tensor, also known as Amari-Chentsov tensor,
$T_{ijk}(m)=\frac{\partial^{3}}{\partial m_{i}\partial m_{j}\partial m_{k}}H(m)=\frac{1}{m_{i}^{2}}\delta_{ij}\delta_{ik},\quad\textrm{for $i,j,k\in\{1,\cdots,d\}$,}$
where $\delta_{ij}$ is a Kronecker delta function. The above Hessian matrix and $3$-tensor are useful in constructing alpha divergences.
Firstly, the Taylor expansion of alpha divergence holds:
$\begin{split}\mathrm{D}_{\alpha}[m\|n]=&\qquad\frac{1}{2}\sum_{i,j=1}^{d}g_{ij}(n)(m_{i}-n_{i})(m_{j}-n_{j})\\ &+\frac{\alpha-3}{12}\sum_{i,j,k=1}^{d}T_{ijk}(n)(m_{i}-n_{i})(m_{j}-n_{j})(m_{k}-n_{k})+O(|m-n|^{4}),\end{split}$
where $|\cdot|$ is the Euclidean norm in $\mathbb{R}^{d}$. Secondly, there are a pair of dual geodesics, namely $\pm\alpha$-geodesics. Denote the $\alpha$-connections at a point $m\in\mathbb{R}_{+}^{d}$ by a three index symbol
$\Gamma^{k,\alpha}_{ij}(m)=-\frac{1+\alpha}{2}m_{i}\cdot T_{ijk}(m).$
Then the alpha geodesic is given below. Denote $\gamma_{\alpha}(t)\in\mathbb{R}_{+}^{d}$, $t\in[0,1]$, with initial point and terminal point $\gamma_{\alpha}(0)=m$, $\gamma_{\alpha}(1)=n$, and
$\frac{d^{2}}{dt^{2}}\gamma_{\alpha}(t)_{k}+\sum_{i,j=1}^{d}\Gamma^{k,\alpha}_{ij}(\gamma_{\alpha}(t))\frac{d}{dt}\gamma_{\alpha}(t)_{i}\frac{d}{dt}\gamma_{\alpha}(t)_{j}=0.$ (2)
Note that the above ODE has a closed-form solution after a change variable, namely $\alpha$-representation
$k_{\alpha}(z)=\left\{\begin{aligned} &\frac{2}{1-\alpha}(z^{\frac{1-\alpha}{2}}-1),&\quad\alpha\neq 1;\\ &\log z,&\quad\alpha=1.\end{aligned}\right$. (3)
Hence $\frac{d^{2}}{dt^{2}}k_{\alpha}(\gamma_{\alpha}(t))=0$. Thus, if $\alpha\neq 1$, the solution of alpha geodesic satisfies
$\gamma_{\alpha}(t)=\Big{(}(1-t)m^{\frac{1-\alpha}{2}}+tn^{\frac{1-\alpha}{2}}\Big{)}^{\frac{2}{1-\alpha}}.$
In above formula, the power and product are componentwise. If $\alpha=-1$, then ODE (2)’s solution is named the mixture (m)-geodesics:
$\gamma_{-1}(t)=(1-t)m+tn.$
If $\alpha=1$, then ODE (2)’s solution is called the exponential (e)-geodesics:
$\gamma_{1}(t)=m^{(1-t)}n^{t}.$
If $\alpha=0$, then (2)’s solution is the Riemannian geodesic of Fisher metric in positive octant:
$\gamma_{0}(t)=\Big{(}(1-t)m^{\frac{1}{2}}+tn^{\frac{1}{2}}\Big{)}^{2}.$
With above defined alpha geodesics, there are duality properties of alpha divergences, including Bregman divergences in terms of $\alpha$-representations (3), and generalized Pythagorean theorem. In literature [2, 26, 31], $(\mathbb{R}_{+}^{d},g,T)$ is an example of $\alpha$-geometry, or Hessian structure of entropy function $H$.
3. Transport alpha divergences
In this section, we define alpha divergences in Wasserstein-$2$ space. Several properties are presented, including composite Bregman divergences and generalized Pythagorean theorem in Wasserstein-$2$ spaces. We also define the alpha geodesic for the completeness of the result.
3.1. Review of Wasserstein-$2$ distances
We briefly review some basic facts on Wasserstein-$2$ distances [3]. We only consider the one dimensional case. In the statistics literatue, the Wasserstein distance will be represented by the quantile functions [12].
Denote a one-dimensional sample space $\Omega=\mathbb{R}^{1}$ with the Euclidean distance. Write a strictly positive probability density space by
$\mathcal{P}(\Omega)=\Big{\{}p\in C(\Omega)\colon\int_{\Omega}p(x)dx=1,~p(x)>0\Big{\}}.$
where $\int$, $dx$ are standard integration symbols in $1D$. For any two continuous probability densities $p$, $q\in\mathcal{P}(\Omega)$ with finite second moments, the Wasserstein-$2$ distance [3, 30] is defined by:
$\textrm{W}_{2}(p,q):=\inf_{T\colon\Omega\rightarrow\Omega}\sqrt{\int_{\Omega}|T(x)-x|^{2}q(x)dx},$ (4)
where the infimum is taken over all continuous mapping function $T$ that pushforwards $q$ to $p$. In other words, $T_{\#}q=p$, which means the Monge-Amperé equation holds:
$p(T(x))\cdot T^{\prime}(x)=q(x).$ (5)
In one-dimensional space, the optimal mapping function $T$ is monotone, which can be solved analytically in terms of quantile functions. From now on, we denote the cumulative distribution functions (CDFs) $F_{p}$, $F_{q}$ of probability density function $p$, $q$, such that
$F_{p}(x)=\int_{-\infty}^{x}p(y)dy,\quad F_{q}(x)=\int_{-\infty}^{x}q(y)dy.$
Denote the quantile functions of probability density $p$, $q$ by
$\begin{split}Q_{p}(u)=&\inf\{x\in\mathbb{R}\colon u\leq F_{p}(x)\}=F_{p}^{-1}(u),\\ Q_{q}(u)=&\inf\{x\in\mathbb{R}\colon u\leq F_{q}(x)\}=F_{q}^{-1}(u).\end{split}$
Note that $F_{p}$ and $F_{q}$ are strictly monotonic increasing functions. We write $F_{p}^{-1}$, $F_{q}^{-1}$ are inverse CDFs of $p$, $q$, respectively. We are ready to solve equation (5). Taking the integration on both sides of equation (5) w.r.t. $x$, we have
$F_{p}(T(x))=F_{q}(x).$
From the inverse function of a CDF, the optimal transport mapping function satisfies
$T(x):=F_{p}^{-1}(F_{q}(x))=Q_{p}(F_{q}(x)).$ (6)
From now on, we always use $T(x)$ to represent the optimal mapping function. Equivalently, the squared Wasserstein-$2$ distance can be formulated as follows.
$\begin{split}\textrm{W}_{2}(p,q)^{2}=&{\int_{\Omega}|Q_{p}(F_{q}(x))-x|^{2}q(x)dx}\\ =&{\int_{0}^{1}|Q_{p}(u)-Q_{q}(u)|^{2}du},\end{split}$
where we apply the change of variable $u=F_{q}(x)\in[0,1]$ in the second equality. In other words, the Wasserstein-$2$ distance in one dimension is the $L^{2}$ distance in quantile functions. In other words, Wasserstein-$2$ geometry is flat since $\mathcal{P}(\Omega)$ is isometric to a convex subset of $L^{2}([0,1])$.
3.2. Transport alpha divergences
We are ready to define transport alpha divergence. Denote a one-parameter function $f_{\mathrm{T},\alpha}\colon\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ by
$f_{\mathrm{T},\alpha}(z)=\left\{\begin{aligned} &\frac{1}{\alpha^{2}}(z^{\alpha}-\alpha\log z-1),&\quad\mathrm{if}\quad\alpha\neq 0;\\ &\frac{1}{2}|\log z|^{2},&\quad\mathrm{if}\quad\alpha=0.\end{aligned}\right$.
Definition 1 (Transport alpha divergence).
Define the functional $\mathrm{D}_{\mathrm{T},\alpha}\colon\mathcal{P}(\Omega)\times\mathcal{P}(\Omega)\rightarrow\mathbb{R}$ by
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q):=\int_{\Omega}f_{\mathrm{T},\alpha}(T^{\prime}(x))q(x)dx=\int_{\Omega}f_{\mathrm{T},\alpha}(\frac{q(x)}{p(T(x))})q(x)dx,$
where $T$ is the monotone mapping function that pushforwards $q$ to $p$, such that $T_{\#}q=p$. We name $\mathrm{D}_{\mathrm{T},\alpha}$ the transport alpha divergence.
We can represent the transport alpha divergence in terms of quantile density functions (QDFs). Denote the QDFs of probability densities $p$ and $q$ below.
$Q_{p}^{\prime}(u)=\frac{d}{du}Q_{p}(u),\quad Q^{\prime}_{q}(u)=\frac{d}{du}Q_{q}(u).$
Proposition 1.
The following equation holds:
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\int_{0}^{1}f_{\mathrm{T},\alpha}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})du.$ (7)
Proof.
Denote a variable $u=F_{q}(x)$, $u\in[0,1]$. Thus, by changing $x$ to $u$ in the following integration,
$\begin{split}\int_{\Omega}f_{\mathrm{T},\alpha}(T^{\prime}(x))q(x)dx=&\int_{\Omega}f_{\mathrm{T},\alpha}(\frac{d}{dx}Q_{p}(F_{q}(x)))q(x)dx\\ =&\int_{\Omega}f_{\mathrm{T},\alpha}\Big{(}\frac{\frac{d}{du}Q_{p}(u)|_{u=F_{q}(x)}}{1/\frac{dF_{q}(x)}{dx}}\Big{)}q(x)dx\\ =&\int_{0}^{1}f_{\mathrm{T},\alpha}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})du,\end{split}$
where the last equality applies the chain rule that $1/\frac{dF_{q}(x)}{dx}=\frac{dx}{dF_{q}(x)}=\frac{d}{du}Q_{q}(u)$. This finishes the proof. ∎
We next present several examples of transport alpha divergences.
  • (i)
    $\alpha=1$: transport KL divergence [17]
    $\mathrm{D}_{\mathrm{T},1}(p\|q)=\int_{0}^{1}\Big{(}\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}-\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}-1\Big{)}du.$
  • (ii)
    $\alpha=-1$: transport reverse KL divergence
    $\mathrm{D}_{\mathrm{T},-1}(p\|q)=\int_{0}^{1}\Big{(}\frac{Q_{q}^{\prime}(u)}{Q_{p}^{\prime}(u)}-\log\frac{Q_{q}^{\prime}(u)}{Q_{p}^{\prime}(u)}-1\Big{)}du.$
  • (iii)
    $\alpha=0$: transport Hessian distance [18] (up to a scaling factor)
    $\mathrm{D}_{\mathrm{T},0}(p\|q)=\frac{1}{2}\int_{0}^{1}|\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}|^{2}du.$
We also present transport alpha divergences with $\alpha=\pm 3$.
  • (iv)
    $\alpha=3$: transport Chi-square divergence
    $\mathrm{D}_{\mathrm{T},3}(p\|q)=\frac{1}{9}\int_{0}^{1}\Big{(}\big{(}\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}\big{)}^{3}-{3}\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}-1\Big{)}du.$
  • (v)
    $\alpha=-3$: transport inverse Chi-square divergence
    $\mathrm{D}_{\mathrm{T},-3}(p\|q)=\frac{1}{9}\int_{0}^{1}\Big{(}\big{(}\frac{Q_{q}^{\prime}(u)}{Q_{p}^{\prime}(u)}\big{)}^{3}-{3}\log\frac{Q_{q}^{\prime}(u)}{Q_{p}^{\prime}(u)}-1\Big{)}du.$
3.3. Properties
In this section, we show that there are several dualities and convexity properties for transport alpha divergences. The proofs are based on the fact that transport alpha divergences are generalized Bregman divergences in Wasserstein-$2$ space.
Proposition 2 (Positivity and Duality).
For any $\alpha\in\mathbb{R}$, and $p$, $q\in\mathcal{P}(\Omega)$, the following properties hold:
  • (i)
    Positivity:
    $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)\geq 0.$
    In addition, $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=0$ if and only if there exists a constant $c\in\mathbb{R}$, such that
    $p(x+c)=q(x).$
  • (ii)
    Duality:
    $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\mathrm{D}_{\mathrm{T},-\alpha}(q|p).$
Proof.
(i) For $\alpha\neq 0$, note that $x-\log x-1\geq 0$ when $x>0$. Thus,
$f_{\mathrm{T},\alpha}(z)=\frac{1}{\alpha^{2}}(z^{\alpha}-\log z^{\alpha}-1)\geq 0.$
Since $q>0$, we have $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)\geq 0$. If $\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=0$, we have $f_{\mathrm{T},\alpha}(T_{p}^{\prime}(x))=0$. Note that $x-\log x-1=0$ iff $x=1$. Thus, $T_{p}^{\prime}(x)=1$. This means that $T(x)=x+c$, where $c$ is a constant. From $(T_{p})_{\#}q=p$, we prove (i) with $\alpha\neq 0$. Similarly, we can prove the result for $\alpha\neq 0$.
(ii) The duality is from equation (7). For any $z_{1}$, $z_{2}>0$, we have $f_{\mathrm{T},\alpha}(\frac{z_{1}}{z_{2}})=f_{\mathrm{T},-\alpha}(\frac{z_{2}}{z_{1}})$. Replacing $z_{1}$, $z_{2}$ by QDFs $Q^{\prime}_{p}$, $Q^{\prime}_{q}$, respectively, and using (7), we finish the proof.
Proposition 3 (Taylor expansions in Wasserstein-$2$ spaces).
The following equation holds:
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{2}\int_{0}^{1}|\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}|^{2}du+\frac{\alpha-3}{6}\int_{0}^{1}\Big{(}\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}\Big{)}^{3}du\\ &+{O(\int_{0}^{1}|\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}|^{4}du)}.\end{split}$
Proof.
We note that
$f_{\mathrm{T},\alpha}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})=f_{\mathrm{T},\alpha}(1+h(u)),$
where we denote a function $h(u):=\frac{Q^{\prime}_{p}(u)-Q^{\prime}_{q}(u)}{Q^{\prime}_{q}(u)}$. By applying a Taylor expansion on $f_{\mathrm{T},\alpha}$ at $1$, we obtain
$f_{\mathrm{T},\alpha}(1+h(u))=f_{\mathrm{T},\alpha}(1)+f^{\prime}_{\mathrm{T},\alpha}(1)h(u)+\frac{1}{2}f_{\mathrm{T},\alpha}^{\prime\prime}(1)|h(u)|^{2}+\frac{1}{6}f_{\mathrm{T},\alpha}^{\prime\prime\prime}(1)h(u)^{3}+O(|h(u)|^{4}).$
Note that $f_{\mathrm{T},\alpha}(1)=f_{\mathrm{T},\alpha}^{\prime}(1)=0$, $f_{\mathrm{T},\alpha}^{\prime\prime}(1)=1$, and $f_{\mathrm{T},\alpha}^{\prime\prime\prime}(1)=\alpha-3$. We finish the proof. ∎
We next represent transport alpha divergences in terms of generalized Bregman divergences in Wasserstein-$2$ spaces. Denote a function $\mathrm{D}_{-1}=\mathrm{D}_{\mathrm{IS}}\colon\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$, such that for $z_{1}$, $z_{2}\in\mathbb{R}_{+}$,
$\mathrm{D}_{-1}(z_{1}|z_{2}):=\frac{z_{1}}{z_{2}}-\log\frac{z_{1}}{z_{2}}-1.$
Here, the notation $\mathrm{D}_{-1}$ is a (1)(-1)-divergence, which is a Bregman divergence with a potential function
$\Psi(z):=-\log z,\quad z\in\mathbb{R}_{+}.$
Here, $\mathrm{D}_{-1}$ is the Itakura–Saito divergence for the point divergence function.
Theorem 1 (Alpha–Itakura–Saito divergences in Wasserstein-$2$ spaces).
Let $\alpha\neq 0$. The following equality holds:
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{\alpha^{2}}\int_{0}^{1}\mathrm{D}_{-1}\big{(}Q_{p}^{\prime}(u)^{\alpha}|Q_{q}^{\prime}(u)^{\alpha}\big{)}du.\end{split}$
In addition, the following generalized Bregman divergence relation holds:
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{\alpha^{2}}\int_{0}^{1}\Big{[}\Psi(Q_{p}^{\prime}(u)^{\alpha})-\Psi(Q^{\prime}_{q}(u)^{\alpha})-\Psi^{\prime}(Q^{\prime}_{q}(u)^{\alpha})\cdot(Q^{\prime}_{p}(u)^{\alpha}-Q^{\prime}_{q}(u)^{\alpha})\Big{]}du.\end{split}$ (8)
Equivalently,
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\quad\frac{1}{\alpha}\Big{[}\int_{\Omega}p(x)\log p(x)dx-\int_{\Omega}q(x)\log q(x)dx\Big{]}\\ &+\frac{1}{\alpha^{2}}\int_{\Omega}\Big{(}(\frac{q(x)}{p(T(x))})^{\alpha}-1\Big{)}q(x)dx.\end{split}$ (9)
Proof.
We first prove equation (8). From equation (7), we have
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=\int_{0}^{1}f_{\mathrm{T},\alpha}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})du=\frac{1}{\alpha^{2}}\int_{0}^{1}{\mathrm{D}_{-1}\big{(}Q_{p}^{\prime}(u)^{\alpha}|Q_{q}^{\prime}(u)^{\alpha}\big{)}}du.$
From the fact of $\mathrm{D}_{-1}$ is a Bregman divergence function, we have
$\mathrm{D}_{-1}(z_{1}|z_{2})=\Psi(z_{1})-\Psi(z_{2})-\Psi^{\prime}(z_{2})\cdot(z_{1}-z_{2}).$
This finishes the proof of (8).
We next prove equation (9). Let $u=F_{q}(x)$, $x=Q_{q}(u)=F_{q}^{-1}(u)$. From the chain rule, we have $\frac{dQ_{q}(u)}{du}=\frac{dx}{dF_{q}(x)}=\frac{1}{q(x)}$, and
$\frac{dQ_{p}(u)}{du}=\frac{dQ_{p}(F_{q}(x))}{dF_{q}(x)}=\frac{\frac{dQ_{p}(F_{q}(x))}{dx}}{\frac{dF_{q}(x)}{dx}}=\frac{T^{\prime}(x)}{q(x)}=\frac{1}{p(T(x))},$ (10)
where the last equality is from the Monge-Amperé equation (5). Let us apply the above estimations to equation (8). We first observe the following fact. Let $u=F_{q}(x)$.
$\begin{split}\int_{0}^{1}\Psi(Q^{\prime}_{q}(u)^{\alpha})du=&-\alpha\int_{0}^{1}\log(\frac{d}{du}Q^{\prime}_{q}(u))du\\ =&-\alpha\int_{0}^{1}\log\frac{1}{q(x)}q(x)dx=\alpha\int_{\Omega}q(x)\log q(x)dx.\end{split}$
Similarly, let $u=F_{p}(x)$, we have
$\int_{0}^{1}\Psi(Q^{\prime}_{p}(u)^{\alpha})du=\alpha\int_{\Omega}p(x)\log p(x)dx.$
We second obtain the following fact. Let $u=F_{q}(x)$, we have
$\begin{split}\int_{0}^{1}\Psi^{\prime}(Q^{\prime}_{q}(u)^{\alpha})\cdot(Q^{\prime}_{p}(u)^{\alpha}-Q^{\prime}_{q}(u)^{\alpha})du=&-\int_{0}^{1}\frac{1}{Q^{\prime}_{q}(u)^{\alpha}}\cdot(Q^{\prime}_{p}(u)^{\alpha}-Q^{\prime}_{q}(u)^{\alpha})du\\ =&-\int_{\Omega}\Big{(}(\frac{q(x)}{p(T(x))})^{\alpha}-1\Big{)}q(x)dx.\end{split}$
Following Theorem 1, we note that the transport alpha divergence is a Bregman divergence in QDFs after a change of variable. We now present the generalized Pythagorean theorem. Denote the Legendre transformation of function $\Psi(z)=-\log z$ below:
$\Psi^{*}(z^{*})=\sup_{z\in\mathbb{R}}\Big{\{}zz^{*}-\Psi(z)\Big{\}}.$
Here $z^{*}=\Psi^{\prime}(z)$, and $\Psi^{*}(z^{*})+\Psi(z)=zz^{*}$. Thus, $z^{*}=-\frac{1}{z}$, and $\Psi^{*}(z^{*})=-\log(-z^{*})-1$.
Corollary 2 (Generalized Pythagorean theorem in Wasserstein-$2$ spaces).
Let $p$, $q$, $r$ be three probability density functions in $\mathcal{P}(\Omega)$. Assume that the following orthogonal condition holds:
$\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\int_{0}^{1}\big{(}Q^{\prime}_{p}(u)^{\alpha}-Q^{\prime}_{q}(u)^{\alpha}\big{)}\cdot\big{(}\frac{1}{Q^{\prime}_{r}(u)^{\alpha}}-\frac{1}{Q^{\prime}_{q}(u)^{\alpha}}\big{)}du=0,&\quad\mathrm{if}\quad\alpha\neq 0;\\ &\int_{0}^{1}\log\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}\cdot\log\frac{Q^{\prime}_{r}(u)}{Q^{\prime}_{q}(u)}du=0,&\quad\mathrm{if}\quad\alpha\neq 0.\end{aligned}\right$. (11)
Then
$\mathrm{D}_{\mathrm{T},\alpha}(p\|q)+\mathrm{D}_{\mathrm{T},\alpha}(q|r)=\mathrm{D}_{\mathrm{T},\alpha}(p|r).$
Proof.
The proof follows from the definition of Bregman divergences. We note the fact that for $z_{1}$, $z_{2}>0$,
$\mathrm{D}_{\mathrm{IS}}(z_{1}|z_{2})=\Psi(z_{1})+\Psi^{*}(z_{2}^{*})-z_{1}\cdot z_{2}^{*}.$
Let $\alpha\neq 0$. Denote $K_{p}=Q^{\prime}_{p}(u)^{\alpha}$ and $K_{p}^{*}=-\frac{1}{Q^{\prime}_{p}(u)^{\alpha}}$, for any $p\in\mathcal{P}(\Omega)$. From equation (8), we have
$\begin{split}&\mathrm{D}_{\mathrm{T},\alpha}(p\|q)+\mathrm{D}_{\mathrm{T},\alpha}(q|r)\\ =&\frac{1}{\alpha^{2}}\int_{0}^{1}\Big{[}\Psi(K_{p})+\Psi^{*}(K_{q}^{*})-K^{*}_{q}\cdot K_{p}+\Psi(K_{q})+\Psi^{*}(K_{r}^{*})-K^{*}_{r}\cdot K_{q}\Big{]}du\\ =&\frac{1}{\alpha^{2}}\int_{0}^{1}\Big{[}\Psi(K_{p})+\Psi^{*}(K_{r}^{*})-K_{p}\cdot K^{*}_{r}+K_{p}\cdot K^{*}_{r}+K_{q}\cdot K_{q}^{*}-K^{*}_{q}\cdot K_{p}-K^{*}_{r}\cdot K_{q}\Big{]}du\\ =&\mathrm{D}_{\mathrm{T},\alpha}(p|r)+\frac{1}{\alpha^{2}}\int_{0}^{1}(K_{p}-K_{q})\cdot(K^{*}_{r}-K^{*}_{q})du.\end{split}$
From the orthogonal condition (11), we finish the proof for $\alpha\neq 0$. For $\alpha=0$, the proof is from the fact that we use the coordinate $\hat{K}_{p}=\log Q^{\prime}_{p}(u)$, under which the transport alpha divergence is an Euclidean distance. The Pythagorean theorem is simple to show. ∎
We can also present the orthogonal condition (11) in terms of pushforward mapping functions.
Corollary 3 (Transport orthogonal condition).
Orthogonal condition (11) is equivalent to
$\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\int_{\Omega}\Big{(}\frac{1}{p(T_{p}(x))^{\alpha}}-\frac{1}{q(x)^{\alpha}}\Big{)}\cdot\Big{(}r(T_{r}(x))^{\alpha}-q(x)^{\alpha}\Big{)}q(x)dx=0,&\quad\mathrm{if}\quad\alpha\neq 0;\\ &\int_{\Omega}\log\frac{q(x)}{p(T_{p}(x))}\cdot\log\frac{q(x)}{r(T_{r}(x))}q(x)dx=0,&\quad\mathrm{if}\quad\alpha=0,\end{aligned}\right$.
where $T_{p}$, $T_{r}$ are monotone functions pusforward $q$ to $p$, $r$, respectively. I.e., $(T_{p})_{\#}q=p$, $(T_{r})_{\#}q=r$.
Proof.
We let $u=F_{q}(x)$. From equation (10), we have
$\frac{dQ_{p}(u)}{du}=\frac{1}{p(T_{p}(x))},\quad\frac{dQ_{r}(u)}{du}=\frac{1}{r(T_{r}(x))}.$
We finish the proof by substituting the above formulas into condition (11). ∎
Remark 1.
We remark that the proposed divergences are canonical divergences [6] associated with the transport Hessian metric. In the appendix, we first review the transprot Hessian meric, the transport KL divergence, and then provide the proof of canonical divergence.
3.4. Transport alpha geodesic
In this section, we construct a one-parameter family of geodesic equations for quantile density functions. We call them transport alpha geodesics. We also present analytical solutions of transport alpha geodesics.
Definition 2 (Transport alpha geodesic equations).
Given two probability density functions $p$, $q\in\mathcal{P}(\Omega)$ and $\alpha\in\mathbb{R}$, the transport alpha geodesic is defined as below. Denote a mapping function $T_{\alpha}\colon[0,1]\times\Omega\rightarrow\Omega$. Consider a one-parameter family of partial differential equations:
$\partial_{tt}\partial_{x}T_{\alpha}(t,x)-(\alpha+1)\frac{(\partial_{t}\partial_{x}T_{\alpha}(t,x))^{2}}{\partial_{x}T_{\alpha}(t,x)}=0,$ (12)
with boundary conditions $T_{\alpha}(0,x)=x$ and $T_{\alpha}(1,\cdot)_{\#}q=p$. Let the curve $r_{\alpha}(t,\cdot)\in\mathcal{P}(\Omega)$, $t\in[0,1]$, then
$r_{\alpha}(t,\cdot)=T_{\alpha}(t,\cdot)_{\#}q,$
is the solution of transport alpha geodesic.
Proposition 4 (Transport alpha geodesics).
Let $T$ be defined in (6). Assume $T^{\prime}(x)\neq 0$ for all $x\in\Omega$. A solution of transport alpha geodesic is given below. The mapping function $T_{\alpha}$ satisfies
$\partial_{x}T_{\alpha}(t,x)=\left\{\begin{aligned} &\Big{(}(1-t)+t(T^{\prime}(x))^{-\alpha}\Big{)}^{-\frac{1}{\alpha}},&\quad\textrm{if $\alpha\neq 0$;}\\ &(T^{\prime}(x))^{t},&\quad\textrm{if $\alpha=0$.}\end{aligned}\right$.
Equivalently, denote $r_{\alpha}(t,\cdot)=T_{\alpha}(t,\cdot)_{\#}q$, and write $Q_{r_{\alpha}}(t,\cdot)$, $\partial_{u}Q_{r_{\alpha}}(t,u)$ as the quantile function, quantile density function of probability density $r_{\alpha}(t,\cdot)$, respectively. Then the transport alpha geodesics in QDFs satisfies
$\partial_{u}Q_{r_{\alpha}}(t,u)=\left\{\begin{aligned} &\Big{(}(1-t)Q^{\prime}_{q}(u)^{-\alpha}+tQ^{\prime}_{p}(u)^{-\alpha}\Big{)}^{-\frac{1}{\alpha}},&\quad\textrm{if $\alpha\neq 0$;}\\ &Q^{\prime}_{p}(u)^{t}Q^{\prime}_{q}(u)^{1-t},&\quad\textrm{if $\alpha=0$.}\end{aligned}\right$.
Proof.
For $\alpha\neq 0$, a simple calculation shows that equation (12) can be reformulated as
$\partial_{tt}(\partial_{x}T_{\alpha}(t,x))^{-\alpha}=0,$
with $T_{\alpha}(0,x)=x$ and $T_{\alpha}(1,x)=T(x)$. Thus, the function $\partial_{x}T_{\alpha}(t,x)$ is uniquely defined, such that
$\begin{split}(\partial_{x}T_{\alpha}(t,x))^{-\alpha}=&t(\partial_{x}T_{\alpha}(1,x))^{-\alpha}+(1-t)(\partial_{x}T_{\alpha}(0,x))^{-\alpha}\\ =&tT^{\prime}(x)^{-\alpha}+(1-t).\end{split}$
This finishes the first part of the proof. By changing the variable $u=F_{q}(x)$, we have
$\partial_{x}T_{\alpha}(t,x)=\partial_{x}F_{r_{\alpha}}^{-1}(F_{q}(x))=Q_{r_{\alpha}}^{\prime}(u)\cdot\frac{du}{dx}=\frac{\partial_{u}Q_{r_{\alpha}}(t,u)}{Q^{\prime}_{q}(u)},$
for any $t\in[0,1]$. Thus,
$\partial_{u}Q_{r_{\alpha}}(t,u)=\Big{(}(1-t)+t(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)})^{-\alpha}\Big{)}^{-\frac{1}{\alpha}}\cdot Q^{\prime}_{q}(u).$
This finishes the second part of proof. Similar derivations also hold for $\alpha=0$. ∎
Proposition 4 can be explained as follows. If $\alpha=-1$, transport-(1)(-1) geodesic also satisfies the geodesic equation in Wasserstein-$2$ space, which is “transportation flat”, meaning that the flatness in the pushforward mapping functions:
$\partial_{x}T_{-1}(t,x)=(1-t)+t\cdot T^{\prime}(x).$ (13)
While, if $\alpha=1$, the transport-$1$ geodesic is an “inverse Jacobi transportation flat” curve. The mapping function pushforwards the density $q$ to $p$ flatly from the following equation:
$\partial_{x}T_{1}(t,x)=\frac{1}{(1-t)+\frac{t}{T^{\prime}(x)}}.$ (14)
If $\alpha=0$, the transport-0 geodesic is a geodesic equation in the transport Hessian metric of negative Boltzmann-Shannon entropy [18, 19]. From now on, we call (13) the $m$-geodesic in Wasserstein-$2$ space, while name (14) the $e$-geodesic in Wasserstein-$2$ space.
Example 1.
Consider two Gaussian distributions $p=\mathcal{N}(0,\Sigma_{p})$, $q=\mathcal{N}(0,\Sigma_{q})$, where $\Sigma_{p}$, $\Sigma_{q}>0$ are covariances of $p$, $q$, respectively. Then $T(x)=\Sigma_{p}^{\frac{1}{2}}\Sigma_{q}^{-\frac{1}{2}}x$. Thus,
$\partial_{x}T_{-1}(t,x)=(1-t)+t\Sigma_{p}^{\frac{1}{2}}\Sigma_{q}^{-\frac{1}{2}},\quad\partial_{x}T_{0}(t,x)=(\Sigma_{p}^{\frac{1}{2}}\Sigma_{q}^{-\frac{1}{2}})^{t},\quad\partial_{x}T_{1}(t,x)=\frac{1}{(1-t)+t\Sigma_{p}^{-\frac{1}{2}}\Sigma_{q}^{\frac{1}{2}}}.$
The following figure demonstrates the above three alpha geodesics.
Figure 1. Three curves for $\partial_{x}T_{\alpha}(t,x)$ with $\Sigma_{p}=1$, $\Sigma_{q}=4$. Red: $\alpha=1$. Black: $\alpha=0$. Blue $\alpha=-1$.
4. Hessian structures of entropy in Wasserstein-$2$ space
In this section, we formulate the Hessian structures in Wasserstein-$2$ space on one-dimensional sample space. In particular, we derive the 3-symmetric tensor from the third order derivatives of Boltzmann–Shannon entropy in Wasserstein-$2$ space.
4.1. Review of Hessian metric in Wasserstein-$2$ space
We briefly recall some facts about the Wasserstein-$2$ metric [30] and the Wasserstein-$2$ Hessian metric [19]. Denote the smooth, strictly positive probability density space by
$\mathcal{P}_{o}(\Omega)=\Big{\{}p\in C^{\infty}(\Omega)\colon\int_{\Omega}p(x)dx=1,~p(x)>0\Big{\}}.$
Denote the tangent space at $p\in\mathcal{P}_{o}(\Omega)$ by
$T_{p}\mathcal{P}_{o}(\Omega)=\Big{\{}\sigma\in C^{\infty}(\Omega)\colon\int_{\Omega}\sigma(x)=0\Big{\}}.$
Write the cotangent space at $p\in\mathcal{P}_{o}(\Omega)$ by
$T_{p}^{*}\mathcal{P}_{o}(\Omega)=C^{\infty}(\Omega)/\mathbb{R}.$
For any constant $c\in\mathbb{R}$, if $\Phi\in T_{p}^{*}\mathcal{P}_{o}(\Omega)$, then $\Phi(x)+c\in T_{p}^{*}\mathcal{P}_{o}(\Omega)$.
Define an inner product $g_{\mathrm{W}}\colon\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\rightarrow\mathbb{R}$ by
$g_{\mathrm{W}}(p)(\sigma_{1},\sigma_{2})=\int_{\Omega}\Phi^{\prime}_{1}(x)\cdot\Phi^{\prime}_{2}(x)p(x)dx,$
where $\sigma_{i}(x)=-\partial_{x}\big{(}p(x)\Phi^{\prime}_{i}(x)\big{)}$, with $\sigma_{i}\in T_{p}\mathcal{P}_{o}(\Omega)$ and $\Phi_{i}\in T^{*}_{p}\mathcal{P}_{o}(\Omega)$, for $i=1,2$. Thus, $(\mathcal{P}(\Omega),g_{\mathrm{W}})$ forms an infinite-dimensional Riemannian manifold in probability density space. In literature, it is often called density manifold [14] or Wasserstein-$2$ space [28].
The Hessian metric in density manifold $(\mathcal{P}(\Omega),g_{\mathrm{W}})$ is defined as follows. Denote the Boltzmann-Shannon entropy by
$\mathcal{H}(p)=-\int_{\Omega}p(x)\log p(x)dx.$
Denote the Hessian operator of negative $\mathcal{H}(p)$ by a two form in $(\mathcal{P}(\Omega),g_{\mathrm{W}})$. In other words, let $g_{\mathrm{H}}=-\textrm{Hess}_{\mathrm{W}}\mathcal{H}\colon\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\rightarrow\mathbb{R}$, then
$g_{\mathrm{H}}(p)(\sigma_{1},\sigma_{2}):=-\textrm{Hess}_{\mathrm{W}}\mathcal{H}(\sigma_{1},\sigma_{2}):=\int_{\Omega}\Phi^{\prime\prime}_{1}(x)\cdot\Phi^{\prime\prime}_{2}(x)p(x)dx,$ (15)
where $\sigma_{i}(x)=-\partial_{x}\big{(}p(x)\Phi^{\prime}_{i}(x)\big{)}$, with $\sigma_{i}\in T_{p}\mathcal{P}_{o}(\Omega)$ and $\Phi_{i}\in T_{p}^{*}\mathcal{P}_{o}(\Omega)$, for $i=1,2$.
4.2. Transport 3–symmetric tensor
We are now ready to formulate the third derivative of entropy $\mathcal{H}(p)$ in Wasserstein-$2$ space. It is a three form, or 3–symmetric tensor in $(\mathcal{P}(\Omega),g_{\mathrm{W}})$.
Definition 3 (Transport 3–symmetric tensor).
Denote $T_{\mathrm{H}}\colon\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\times T_{p}\mathcal{P}_{o}(\Omega)\rightarrow\mathbb{R}$. Then
$T_{\mathrm{H}}(p)(\sigma_{1},\sigma_{2},\sigma_{3})=2\int_{\Omega}\Phi^{\prime\prime}_{1}(x)\cdot\Phi^{\prime\prime}_{2}(x)\cdot\Phi^{\prime\prime}_{3}(x)p(x)dx,$
where $\sigma_{i}(x)=-\big{(}p(x)\Phi^{\prime}_{i}(x)\big{)}^{\prime}$, with $\sigma\in T_{p}\mathcal{P}_{o}(\Omega)$, and $\Phi_{i}\in T^{*}_{p}\mathcal{P}_{o}(\Omega)$, for $i=1,2,3$.
We also present that the transport 3-symmetric tensor introduces a third-order iterative Bakry–Émery Gamma calculus; see related studies between Gamma calculuses [8] and optimal transport [22, 25].
Theorem 4 (Gamma calculus induced 3–symmetric tensor).
Denote bilinear forms $\Gamma_{1}$, $\Gamma_{2}\colon C^{\infty}(\Omega)\times C^{\infty}(\Omega)\rightarrow C^{\infty}(\Omega)$ by
$\Gamma_{1}(\Phi,\Phi)(x)=\Phi^{\prime}(x)\cdot\Phi^{\prime}(x),\quad\Gamma_{2}(\Phi,\Phi)(x)=\Phi^{\prime\prime}(x)\cdot\Phi^{\prime\prime}(x).$
Define the Gamma-3 operator $\Gamma_{3}\colon C^{\infty}(\Omega)\times C^{\infty}(\Omega)\times C^{\infty}(\Omega)\rightarrow C^{\infty}(\Omega)$ by
$\Gamma_{3}(\Phi,\Phi,\Phi)(x):=\Gamma_{2}(\Gamma_{1}(\Phi,\Phi),\Phi)(x)-\Gamma_{1}(\Gamma_{2}(\Phi,\Phi),\Phi)(x).$
Then the following equation holds:
$T_{\mathrm{H}}(p)(\sigma,\sigma,\sigma)=\int_{\Omega}\Gamma_{3}(\Phi(x),\Phi(x),\Phi(x))p(x)dx,$
where $\sigma=-\partial_{x}(p(x)\Phi^{\prime}(x))$.
Proof.
The proof follows by a direct calculation. Note that
$\Gamma_{1}(\Gamma_{2}(\Phi,\Phi),\Phi)=\partial_{x}(|\Phi^{\prime\prime}|^{2})\Phi^{\prime}=2\Phi^{\prime\prime\prime}\cdot\Phi^{\prime\prime}\cdot\Phi^{\prime},$
and
$\Gamma_{2}(\Gamma_{1}(\Phi,\Phi),\Phi)=\partial_{x}^{2}(|\Phi^{\prime}|^{2})\Phi^{\prime\prime}=2\Phi^{\prime\prime\prime}\cdot\Phi^{\prime\prime}\cdot\Phi^{\prime}+2|\Phi^{\prime\prime}|^{3}.$
By taking the difference between the two functionals, we derive the result. ∎
We finish this section by representing the Taylor expansions of transport alpha divergences, using Hessian structures $(\mathcal{P}_{o}(\Omega),g_{\mathrm{H}},T_{\mathrm{H}})$.
Corollary 5 (Taylor expansions in transport Hessian structures).
For any $p$, $q\in\mathcal{P}_{o}(\Omega)$. Denote $\Phi\in T^{*}_{q}\mathcal{P}_{o}(\Omega)$, such that
$\Phi(x)=\int_{0}^{x}Q_{p}(F_{q}(y))dy-\frac{x^{2}}{2}+c,$
where $c\in\mathbb{R}$ is a constant. Denote $\sigma=\partial_{x}(q(x)\Phi^{\prime}(x))\in T_{q}\mathcal{P}_{o}(\Omega)$. Then, the following equation holds.
$\begin{split}\mathrm{D}_{\mathrm{T},\alpha}(p\|q)=&\frac{1}{2}g_{\mathrm{H}}(q)(\sigma,\sigma)+\frac{\alpha-3}{6}T_{\mathrm{H}}(q)(\sigma,\sigma,\sigma)+{O(\int_{\Omega}|\Phi^{\prime\prime}(x)|^{4}q(x)dx)}.\end{split}$
Proof.
The proof is based on a direct calculation. Note that
$\Phi^{\prime}(x)=Q_{p}(F_{q}(x))-x,$
and
$\Phi^{\prime\prime}(x)=\frac{d}{dx}Q_{p}(F_{q}(x))-1=\frac{Q^{\prime}_{p}(F_{q}(x))}{\frac{1}{q(x)}}-1.$
For $k=2,3$, from the change of variable $u=F_{q}(x)$, we have
$\int_{\Omega}(\Phi^{\prime\prime}(x))^{k}p(x)dx=\int_{0}^{1}(\frac{Q^{\prime}_{p}(u)}{Q^{\prime}_{q}(u)}-1)^{k}du.$
From Proposition 3, we finish the proof. ∎
Remark 2.
We note that $\Gamma_{1}$, $\Gamma_{2}$ are often called Gamma one and Gamma two operators, which are firstly introduced by Bakry–Émery [8] to study the Ricci curvature lower bound on a sample space. For simplicity of presentation, we only show them in one-dimensional sample space. The iterative Gamma two calculus connects with second-order derivatives of entropy in Wasserstein-$2$ space [8, 30] with generalizations [15]. Here, we present a “third-order” Gamma calculus to formulate the third derivatives in Wasserstein-$2$ space, namely transport 3-symmetric tensor. We will study geometric calculations of transport-3 symmetric tensors in high-dimensional spaces in future works. Following [15, 16], we expect that information geometry and Gamma three operators are tools in studying generalized divergences in Wasserstein-$2$ type spaces.
5. Examples
This section provides examples of transport alpha divergences between one-dimensional probability distributions, including generative models, location-scale families, and Cauchy distributions.
In machine learning applications [4], a generative model is defined as follows. Consider a latent random variable $Z\sim p_{\mathrm{ref}}$, where $p_{\mathrm{ref}}\in\mathcal{P}(\Omega)$ is a given reference measure. Denote an invertible map function $G\colon\Omega\times\Theta\rightarrow\Omega$, where $\Theta\subset\mathbb{R}^{n}$ is a parameter space. Then
$G(\cdot,\theta)_{\#}p_{\textrm{ref}}(\cdot)=p(\cdot,\theta).$
If $G$ is linear w.r.t. $Z$, the generative family forms a location-scale family. Furthermore, if $G$ is linear and $Z$ follows a Gaussian distribution, the generative model formulates a class of Gaussian distributions.
Proposition 5 (Transport alpha divergence in generative models).
Let $\theta_{X}$, $\theta_{Y}\in\Theta$ and consider $Z\sim p_{\mathrm{ref}}$, with
$X=G(Z,\theta_{X})\sim p_{X},\quad Y=G(Z,\theta_{Y})\sim p_{Y}.$
Then the transport alpha divergence between probability distributions $p_{X}$, $p_{Y}$ satisfies
$\mathrm{D}_{\mathrm{T},\alpha}(p_{X}\|p_{Y})=\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\mathbb{E}_{Z\sim p_{\mathrm{ref}}}\Big{[}\Big{(}\frac{\partial_{Z}G(Z,\theta_{X})}{\partial_{Z}G(Z,\theta_{Y})}\Big{)}^{\alpha}-\alpha\log\frac{\partial_{Z}G(Z,\theta_{X})}{\partial_{Z}G(Z,\theta_{Y})}-1\Big{]},&\quad\textrm{if $\alpha\neq 0$;}\\ &\frac{1}{2}\mathbb{E}_{Z\sim p_{\mathrm{ref}}}\Big{[}\Big{(}\log\frac{\partial_{Z}G(Z,\theta_{X})}{\partial_{Z}G(Z,\theta_{Y})}\Big{)}^{2}\Big{]},&\quad\textrm{if $\alpha=0$.}\end{aligned}\right$.
Here $\mathbb{E}$ is the expectation operator. We also compare transport alpha divergences with the Wasserstein-$2$ distance
$W_{2}(p,q)=\sqrt{\mathbb{E}_{Z\sim p_{\mathrm{ref}}}\Big{[}\big{|}G(Z,\theta_{X})-G(Z,\theta_{Y})\big{|}^{2}\Big{]}},$
where we need to assume that $\mathbb{E}_{Z\sim p_{\mathrm{ref}}}|G(Z,\theta)|^{2}<+\infty$, for $\theta=\theta_{X}$ or $\theta_{Y}$.
Example 2 (Location scale family).
Suppose $G$ is a linear mapping function such that
$G(Z,\theta)=\theta Z,$
with $\theta>0$ and $Z\in\mathbb{R}^{1}$. Then $p(\cdot,\theta)=G(\cdot,\theta)\#p_{\textrm{ref}}$ is a location scale family. In this case, we have
$\mathrm{D}_{\mathrm{T},\alpha}(p_{X}\|p_{Y})=\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\Big{[}\Big{(}\frac{\theta_{X}}{\theta_{Y}}\Big{)}^{\alpha}-\alpha\log\frac{\theta_{X}}{\theta_{Y}}-1\Big{]},&\quad\textrm{if $\alpha\neq 0$;}\\ &\frac{1}{2}\Big{(}\log\frac{\theta_{X}}{\theta_{Y}}\Big{)}^{2},&\quad\textrm{if $\alpha=0$.}\end{aligned}\right$.
We last present an example of the Wasserstein-$2$ distance not being well defined, meaning that the distributions are not with the finite second moment. In this case, the transport alpha divergence is still well defined.
Example 3 (Cauchy distributions).
The Cauchy distribution is defined as follows. For $\gamma>0$,
$p(x,\gamma)=\frac{1}{\pi\gamma}\Big{[}\frac{1}{(\frac{x}{\gamma})^{2}+1}\Big{]}.$
Thus, denote $T(x)=\gamma\cdot x$, we have $T_{\#}p(\cdot,1)=p(\cdot,\gamma)$. For $\gamma_{1}$, $\gamma_{2}>0$, we have
$\mathrm{D}_{\mathrm{T},\alpha}(p(\cdot,\gamma_{1})\|p(\cdot,\gamma_{2}))=\left\{\begin{aligned} &\frac{1}{\alpha^{2}}\Big{[}\Big{(}\frac{\gamma_{1}}{\gamma_{2}}\Big{)}^{\alpha}-\alpha\log\frac{\gamma_{1}}{\gamma_{2}}-1\Big{]},&\quad\textrm{if $\alpha\neq 0$;}\\ &\frac{1}{2}\Big{(}\log\frac{\gamma_{1}}{\gamma_{2}}\Big{)}^{2},&\quad\textrm{if $\alpha=0$.}\end{aligned}\right$.
We remark that the Cauchy distribution does not have the finite second moment. And the Wasserstein-$2$ distance is not well defined, such that $W_{2}(p(\cdot,\gamma_{1}),p(\cdot,\gamma_{2}))=+\infty$.
6. Discussion
This paper proposes a class of transport alpha divergences, one-parameter variation of transport KL divergence and transport Hessian distance. They are connected with Hessian metrics and 33–symmetric tensors of the negative Boltzmann-Shannon entropy in Wasserstein-$2$ space. We provide several analytical examples in one-dimensional probability densities, including generative models.
It is worth mentioning that the quantile density functions (QDFs) have been applied in statistical learning problems [29]. The quantile density functions measure densities’ shape up to any constant shifting. The transport alpha divergence provides a class of functionals for measuring the difference from QDFs, i.e., Jacobi functions of mapping functions. In future work, we shall study transport alpha divergences in high dimensional probability densities [10, 19]. This direction includes analysis, dualities, invariance properties, and optimization algorithms of transport mapping-related divergence functionals. In particular, systemic geometric calculations for Hessian structures in Wasserstein-$2$ spaces $(\mathcal{P}_{o}(\Omega),g_{\mathrm{H}},T_{\mathrm{H}})$ will be investigated; see related studies in [9, 19, 20, 26]. The systemic understanding of three symmetric tensors and alpha connections rely on both studies in optimal transport and information geometry, which is not a simple direction; see related studies in [5, 6, 24]. We expect that the convexity analysis and approximations in transport Hessian structures serve the mathematical foundations of artificial intelligence, in particular for studying new optimization problems or approximations in generative models.
Acknowledgements. W. Li’s work is supported by AFOSR YIP award No. FA9550-23-1-0087, NSF RTG: 2038080, and NSF DMS: 2245097.
Appendix
In this section, we present some necessary proofs for this paper.
6.1. Derivatives in Wasserstein-$2$ space
We first present first, second, and third-order derivatives in Wasserstein-$2$ space. This provides the derivation for transport 33–symmetric tensor defined in Definition 3.
Proposition 6.
Denote $p\colon[0,1]\times\Omega\rightarrow\mathbb{R}$ satisfying the geodesics equation in $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$ with $p(0,x)=p(x)$, $\partial_{t}p(0,x)=\sigma(x)=-\partial_{x}(p(x)\Phi^{\prime}(x))$. Then
$\frac{d^{n}}{dt^{n}}\mathcal{H}(p(t,\cdot))=(-1)^{n}(n-1)!\int_{\Omega}(\Phi^{\prime\prime}(x))^{n}p(x)dx.$
In particular, for $n=1,2,3$, we have
  • (i)
    $\frac{d}{dt}\mathcal{H}(p(t,\cdot))|_{t=0}=\mathrm{grad}_{\mathrm{W}}\mathcal{H}(p)(\sigma)=\int_{\Omega}\Phi^{\prime\prime}(x)p(x)dx.$
  • (ii)
    $\frac{d^{2}}{dt^{2}}\mathcal{H}(p(t,\cdot))|_{t=0}=\mathrm{Hess}_{\mathrm{W}}\mathcal{H}(p)(\sigma,\sigma)=\int_{\Omega}(\Phi^{\prime\prime}(x))^{2}p(x)dx.$
  • (iii)
    $\frac{d^{3}}{dt^{3}}\mathcal{H}(p(t,\cdot))|_{t=0}=T_{\mathrm{H}}(p)(\sigma,\sigma,\sigma)=2\int_{\Omega}(\Phi^{\prime\prime}(x))^{3}p(x)dx.$
Proof.
We recall that the geodesics in $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$ satisfies
$\left\{\begin{split}&\partial_{t}p(t,x)+\partial_{x}(p(t,x)\partial_{x}\Phi(t,x))=0\\ &\partial_{t}\Phi(t,x)+\frac{1}{2}|\partial_{x}\Phi(t,x)|^{2}=0,\end{split}\right$.
where $p(0,x)=p(x)$ and $\partial_{t}p(0,x)=\sigma(x)=-\partial_{x}(p(x)\Phi^{\prime}(x))$. We prove the result by induction. When $n=1$, we have
$\begin{split}-\frac{d}{dt}\mathcal{H}(p(t,\cdot))|_{t=0}=&-\int_{\Omega}\partial_{x}(p(x)\Phi^{\prime}(x))(\log p(x)+1)dx\\ =&\int_{\Omega}\Phi^{\prime}(x)\partial_{x}\log p(x)p(x)dx\\ =&\int_{\Omega}\Phi^{\prime}(x)\partial_{x}p(x)dx\\ =&-\int_{\Omega}\Phi^{\prime\prime}(x)p(x)dx,\end{split}$
where we use the fact that $\partial_{x}\log p(x)\cdot p(x)=\frac{\partial_{x}p(x)}{p(x)}\cdot p(x)=\partial_{x}p(x)$ in the third equality. Assume that for $n=k$, $k\in\mathbb{N}$, we have
$\frac{d^{k}}{dt^{k}}\mathcal{H}(p(t,\cdot))|_{t=0}=(-1)^{k}(k-1)!\int_{\Omega}(\Phi^{\prime\prime}(x))^{k}p(x)dx.$
Note that the second equation of the geodesic in $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$ can be reformulated as below:
$\partial_{t}\partial_{x}\Phi(t,x)+\partial_{xx}\Phi(t,x)\cdot\partial_{x}\Phi(t,x)=0.$
Hence
$\begin{split}\frac{d}{dt}\int_{\Omega}(\partial_{xx}\Phi(t,x))^{k}p(t,x)dx=&\int_{\Omega}\partial_{t}\Big{(}(\partial_{xx}\Phi(t,x))^{k}\Big{)}p(t,x)dx+\int_{\Omega}(\partial_{xx}\Phi(t,x))^{k}\partial_{t}p(t,x)dx\\ =&\quad\int_{\Omega}k(\partial_{xx}\Phi(t,x))^{k-1}\partial_{x}^{2}\partial_{t}\Phi(t,x)p(t,x)dx\\ &-\int_{\Omega}(\partial_{xx}\Phi(t,x))^{k}\partial_{x}(p(t,x)\partial_{x}\Phi(t,x))dx\\ =&-\int_{\Omega}k(\partial_{xx}\Phi(t,x))^{k-1}\partial_{x}(\partial_{xx}\Phi(t,x)\partial_{x}\Phi(t,x))p(t,x)dx\\ &+\int_{\Omega}k(\partial_{xx}\Phi(t,x))^{k-1}\partial_{x}^{3}\Phi(t,x)\partial_{x}\Phi(t,x)p(t,x)dx\\ =&-\int_{\Omega}k(\partial_{xx}\Phi(t,x))^{k+1}p(t,x)dx.\end{split}$
From the assumption, we have
$\begin{split}-\frac{d^{k+1}}{dt^{k+1}}\mathcal{H}(p(t,\cdot))|_{t=0}=&(-1)^{k}(k-1)!\frac{d}{dt}\int_{\Omega}(\Phi^{\prime\prime}(t,x))^{k}p(t,x)dx|_{t=0}\\ =&(-1)^{k}(k-1)!\cdot(-1)\cdot k\int_{\Omega}(\Phi^{\prime\prime}(x))^{k+1}p(x)dx\\ =&(-1)^{k+1}k!\int_{\Omega}(\Phi^{\prime\prime}(x))^{k+1}p(x)dx,\end{split}$
which finishes the proof. ∎
Remark 3.
These geometric formulas are derived based on the Riemannian Levi-Civita connection in density manifold $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$. They formulate classical Gamma calculuses; see details in [15, 16, 30]. We leave studies of high-order derivatives of entropy in $(\mathcal{P}_{o}(\Omega),g_{\mathrm{W}})$ with high dimensional sample spaces in future works.
6.2. Transport Hessian distance
We next review the definition of transport Hessian distances between probability distributions. For completeness of this paper, we also provide its derivation here.
Define a distance function $\mathrm{Dist}_{\mathrm{H}}$ $\colon\mathcal{P}(\Omega)\times\mathcal{P}(\Omega)\rightarrow\mathbb{R}$ by
$\mathrm{Dist}_{\mathrm{H}}(p,q)^{2}=\inf_{p\colon[0,1]\times\Omega\rightarrow\mathbb{R}}\Big{\{}\int_{0}^{1}g_{\mathrm{H}}(\partial_{t}p,\partial_{t}p)dt\colon p(0,x)=q(x),~p(1,x)=p(x)\Big{\}}.$ (16)
Here the infimum is taken among all smooth density paths $p\colon[0,1]\times\Omega\rightarrow\mathbb{R}$, which connects both initial and terminal time probability density functions $q$, $p\in\mathcal{P}(\Omega)$. Interestingly, there is a closed-form update for the variational problem (16).
Proposition 7 ([18]).
. The squared transport Hessian distance has the following formulation.
$\mathrm{Dist}_{\mathrm{H}}(p,q)^{2}=\int_{0}^{1}|\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}|^{2}du=2\mathrm{D}_{\mathrm{T},0}(p\|q).$
Proof.
There are two change of variables to derive the Hessian distance. Firstly, denote $p^{0}(x)=q(x)$ and $p^{1}(x)=p(x)$. Denote the variational problem (16) by
$\begin{split}\mathrm{Dist}_{\mathrm{H}}(p^{0},p^{1})^{2}=\inf_{\Phi,p\colon[0,1]\times\Omega\rightarrow\mathbb{R}}&\Big{\{}\int_{0}^{1}\int_{\Omega}|\partial^{2}_{yy}\Phi(t,y)|^{2}p(t,y)dydt\colon\\ &\hskip 28.45274pt\partial_{t}p(t,y)+\partial_{y}(p(t,y)\partial_{y}\Phi(t,y))=0,~\textrm{fixed $p^{0}$, $p^{1}$}\Big{\}},\end{split}$
where the infimum is among all smooth density paths $p\colon[0,1]\times\Omega\rightarrow\mathbb{R}$ satisfying the continuity equation with the gradient potential vector field. The potential is given by $\Phi\colon[0,1]\times\Omega\rightarrow\mathbb{R}$. Denote
$y=T(t,x),\qquad\partial_{t}T(t,x)=\partial_{y}\Phi(t,y).$
Hence
$\begin{split}\mathrm{Dist}_{\mathrm{H}}(p^{0},p^{1})^{2}=&\inf_{T\colon[0,1]\times\Omega\rightarrow\Omega}\Big{\{}\int_{0}^{1}\int_{\Omega}|\partial_{y}v(t,T(t,x))|^{2}p(t,T(t,x))dT(t,x)dt\colon\\ &\hskip 85.35826ptT(t,\cdot)_{\#}p(0,\cdot)=p(t,\cdot)\Big{\}},\end{split}$
where the infimum is taken among all smooth mapping functions $T\colon[0,1]\times\Omega\rightarrow\Omega$ with $T(0,x)=x$ and $T(1,x)=T(x)$. Thus,
$\begin{split}&\int_{0}^{1}\int_{\Omega}|\partial_{y}\partial_{t}T(t,x)|^{2}p(t,T(t,x))\partial_{x}T(t,x)dxdt\\ =&\int_{0}^{1}\int_{\Omega}|\partial_{x}\partial_{t}T(t,x)\frac{dx}{dy}|^{2}p(t,T(t,x))\partial_{x}T(t,x)dxdt\\ =&\int_{0}^{1}\int_{\Omega}|\partial_{t}\partial_{x}T(t,x)\frac{1}{\partial_{x}T(t,x)}|^{2}{q(x)}dxdt\\ =&\int_{0}^{1}\int_{\Omega}|\partial_{t}\log\partial_{x}T(t,x)|^{2}{q(x)}dxdt,\end{split}$ (17)
where we use the fact that $\partial_{t}\log\partial_{x}T(t,x)=\partial_{t}\partial_{x}T(t,x)\frac{1}{\partial_{x}T(t,x)}$.
Secondly, denote $y=F_{q}(x)$, where $y\in[0,1]$. Again, by using a chain rule for $T(t,\cdot)_{\#}q=p_{t}$ with $p_{t}:=p(t,x)$, we have
$q(x)=\frac{dy}{dx}=\frac{1}{\frac{dx}{dy}}=\frac{1}{Q_{q}^{\prime}(y)},\quad\partial_{x}T(t,x)=\partial_{x}F_{p_{t}}^{-1}(F_{q}(x))=Q_{p_{t}}^{\prime}(y)\frac{dy}{dx}=\frac{Q^{\prime}_{p_{t}}(y)}{Q^{\prime}_{q}(y)}.$ (18)
Under the above change of variables and the fact that $\partial_{t}\log\partial_{x}T(t,x)=\partial_{t}\log Q^{\prime}_{p_{t}}(y)$, we observe that the variation problem (17) satisfies
$\begin{split}\mathrm{Dist}_{\mathrm{H}}(p^{0},p^{1})^{2}=&\inf_{Q^{\prime}_{p_{t}}\colon[0,1]^{2}\rightarrow\mathbb{R}}\Big{\{}\int_{0}^{1}\int_{0}^{1}|\partial_{t}\log Q^{\prime}_{p_{t}}(y)|^{2}dydt\Big{\}},\end{split}$ (19)
where the infimum is taken among all paths $Q^{\prime}_{p_{t}}\colon[0,1]^{2}\rightarrow\mathbb{R}$ with fixed initial and terminal time conditions. By using the Euler-Lagrange equation for variable $Q^{\prime}_{p_{t}}$, we show that the geodesic equation in transport Hessian metric satisfies
$\partial_{tt}\log Q^{\prime}_{p_{t}}(y)=0.$
This means
$\log(Q_{p_{t}}^{\prime}(y))=t\log(Q_{p}^{\prime}(y))+(1-t)\log(Q_{q}^{\prime}(y)),$
and
$\partial_{t}\log(Q_{p_{t}}^{\prime}(y))=\log(Q_{p}^{\prime}(y))-\log(Q_{q}^{\prime}(y)).$
Combining the above facts in the variational problem (17), we finish the proof. ∎
6.3. Transport KL divergence
We also briefly present the motivation of transport KL divergence. Consider a functional $\mathcal{F}\colon\mathcal{P}\rightarrow\mathbb{R}$. We define a class of Bregman divergences in Wasserstein-2 space
$\mathrm{D}_{\mathrm{T},\mathcal{F}}(p\|q)=\mathcal{F}(p)-\mathcal{F}(q)-\int_{\Omega}\Big{(}\partial_{x}\frac{\delta}{\delta q(x)}\mathcal{F}(q),T(x)-x\Big{)}q(x)dx,$
where $p$, $q\in\mathcal{P}(\Omega)$, $\frac{\delta}{\delta q(x)}$ is the $L^{2}$ first variation w.r.t. $q(x)$, and $T_{\#}q=p$. We refer $\mathrm{D}_{\mathrm{T},\mathcal{F}}$ as the transport Bregman divergence. If $\mathcal{F}$ is a second moment functional, i.e. $\mathcal{F}(p)=\int_{\Omega}|x|^{2}p(x)dx$, then $\mathrm{D}_{\mathrm{T},\mathcal{F}}$ forms the Wasserstein-2 distance. If $\mathcal{F}(p)=-\mathcal{H}(p)=\int_{\Omega}p(x)\log p(x)dx$, then
$\mathcal{F}(p)=-\mathcal{H}(p)=\int_{\Omega}\log\frac{q(x)}{T^{\prime}(x)}q(x)dx.$
Thus, the transport Bregman divergence of $\mathcal{F}$ satisfies
$\begin{split}\mathrm{D}_{\mathrm{TKL}}(p\|q)=&\mathrm{D}_{\mathrm{T},\mathcal{F}}(p\|q)\\ =&\int_{\Omega}\log\frac{q(x)}{T^{\prime}(x)}q(x)-q(x)\log q(x)-(\partial_{x}\log q(x),T(x)-x)q(x)dx\\ =&\int_{\Omega}\Big{[}-\log T^{\prime}(x)q(x)-(\partial_{x}q(x),T(x)-x)\Big{]}dx\\ =&\int_{\Omega}\Big{(}T^{\prime}(x)-\log T^{\prime}(x)-1\Big{)}q(x)dx,\end{split}$
where we use the fact that $q(x)\partial_{x}\log q(x)=\partial_{x}q(x)$ in the third equality, and apply the integration by parts in the last equality. From now on, we name $\mathrm{D}_{\mathrm{TKL}}$ the transport KL divergence. Again, by the change of variable (18), we note that $\mathrm{D}_{\mathrm{TKL}}$ can be formulate in terms of quantile density functions:
$\mathrm{D}_{\mathrm{TKL}}(p\|q):=\int_{0}^{1}\Big{(}\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}-\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}-1\Big{)}du=\mathrm{D}_{\mathrm{T},1}(p\|q).$
From this formulation, we observe that $\mathrm{D}_{\mathrm{TKL}}$ is an Itakura–Saito type divergence in term of mapping functions or quantile density functions.
6.4. Validation of canonical divergences
We are now ready to valid that the proposed transport alpha divergence is a canonical divergence, which follows the Definition 4.8 in [7]. In other words, the following proposition holds.
Proposition 8.
Let $r(t,\cdot)$, $t\in[0,1]$, be the probability density, satisfying transport alpha geodesic in Proposition 4, with $\alpha=-1$, $r(0,x)=p(x)$ and $r(1,x)=q(x)$. Then the following equality holds:
$\mathrm{D}_{\mathrm{T},1}(p\|q)=\int_{0}^{1}t\cdot g_{\mathrm{H}}(\partial_{t}r,\partial_{t}r)dt,$
where $g_{\mathrm{H}}$ is the transport Hessian metric defined in (15).
Proof.
We follow the idea of proof in [7], with using the change of variable (18) and transport Hessian metric $g_{\mathrm{H}}$. We denote a monotone function $\hat{T}(x)=T^{-1}(x)$, such that $(T^{-1})_{\#}p=q$. Denote $\hat{T}\colon[0,1]\times\Omega\rightarrow\Omega$ satisfying Proposition 4 with $\alpha=-1$, such that
$\partial_{t}(\partial_{x}\hat{T}(t,x))=\partial_{t}(t\hat{T}^{\prime}(x)+(1-t))=\hat{T}^{\prime}(x)-1.$
We prove the following claim.
Claim:
$\begin{split}-\frac{d^{2}}{dt^{2}}\mathcal{H}(r(t,\cdot))=&g_{\mathrm{H}}(\partial_{t}r,\partial_{t}r).\end{split}$
Proof of Claim.
Similar as the derivation of equation (17), we note that
$\begin{split}g_{\mathrm{H}}(\partial_{t}r,\partial_{t}r)=&\int_{\Omega}|\partial_{t}\partial_{x}\hat{T}(t,x)\frac{1}{\partial_{x}\hat{T}(t,x)}|^{2}{q(x)}dx.\end{split}$
We also note that $\mathcal{H}(r(t,\cdot))=\int_{\Omega}r(t,x)\log r(t,x)dx=\int_{\Omega}p(x)\log\frac{p(x)}{\partial_{x}\hat{T}(t,x)}dx$. Hence
$\begin{split}-\frac{d^{2}}{dt^{2}}\mathcal{H}(r(t,\cdot))=&\int_{\Omega}\frac{d^{2}}{dt^{2}}\log\frac{p(x)}{\partial_{x}\hat{T}(t,x)}p(x)dx\\ =&\int_{\Omega}\Big{(}|\partial_{t}\partial_{x}\hat{T}(t,x)\frac{1}{\partial_{x}\hat{T}(t,x)}|^{2}-\frac{\partial_{tt}\partial_{x}\hat{T}(t,x)}{\partial_{x}\hat{T}(t,x)}\Big{)}{q(x)}dx\\ =&g_{\mathrm{H}}(\partial_{t}r,\partial_{t}r).\end{split}$
Here we use the fact that $\partial^{2}_{tt}\partial_{x}\hat{T}(t,x)=0$. ∎
Thus,
$\begin{split}&\int_{0}^{1}t\cdot g_{\mathrm{H}}(\partial_{t}r,\partial_{t}r)dt=\int_{0}^{1}t\cdot(-\frac{d^{2}}{dt^{2}}\mathcal{H}(r(t,\cdot)))dt\\ =&t\cdot(-\frac{d}{dt}\mathcal{H}(r)|_{t=0}^{t=1})-\int_{0}^{1}(-\frac{d}{dt}\mathcal{H}(r))dt\\ =&-\frac{d}{dt}\mathcal{H}(r)|_{t=1}+\mathcal{H}(q)-\mathcal{H}(p).\end{split}$
Here we note that
$\begin{split}-\frac{d}{dt}\mathcal{H}(r)|_{t=1}=&\int_{\Omega}p(x)\frac{d}{dt}\log\frac{p(x)}{\hat{T}^{\prime}(x)}dx|_{t=1}\\ =&-\int_{\Omega}p(x)\partial_{t}\log\hat{T}^{\prime}(t,x)dx|_{t=1}\\ =&-\int_{\Omega}p(x)\frac{\partial_{t}\hat{T}^{\prime}(t,x)}{\hat{T}^{\prime}(t,x)}dx|_{t=1}\\ =&-\int_{\Omega}p(x)\frac{\hat{T}^{\prime}(x)-1}{\hat{T}^{\prime}(x)}dx=-1+\int_{\Omega}\frac{1}{\hat{T}^{\prime}(x)}p(x)dx.\end{split}$ (20)
Similar as in the change of variable formula (18), we have
$\int_{\Omega}\frac{1}{\hat{T}^{\prime}(x)}p(x)dx=\int_{0}^{1}\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}du.$
We also note the fact that
$\begin{split}\mathcal{H}(q)-\mathcal{H}(p)=&\int_{\Omega}\Big{[}p(x)\log p(x)-q(x)\log q(x)\Big{]}dx\\ =&\int_{\Omega}p(x)\log p(x)dx-\int_{\Omega}\log q(\hat{T}(x))q(\hat{T}(x))\hat{T}^{\prime}(x)dx\\ =&\int_{\Omega}p(x)\log p(x)dx-\int_{\Omega}\log\frac{p(x)}{\hat{T}^{\prime}(x)}p(x)dx\\ =&\int_{\Omega}p(x)\log\hat{T}^{\prime}(x)dx\\ =&-\int_{0}^{1}\log\frac{Q_{p}^{\prime}(u)}{Q_{q}^{\prime}(u)}du,\end{split}$ (21)
where the last equality again follows from the change of variable formula in (18). Combining the above equalities in (20) and (21), we finish the proof. ∎